OpenStudy (anonymous):

What is the volume of a square prism whose base edges are 2 centimeters and height is 8 centimeters? 128 cm3 8 cm3 16 cm3 32 cm3

OpenStudy (anonymous):

VOLUME OF A PRISM = area of base * height

OpenStudy (anonymous):

now can u?

OpenStudy (anonymous):

use de formula fr volume of square prism.. lenght*breadth*height

OpenStudy (anonymous):

hey ronaldo

OpenStudy (anonymous):

the anwser is 16in??? rite??? cuz itz area of base*height?????

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

at first calculate area of base and tell me

OpenStudy (anonymous):

itz 8

OpenStudy (anonymous):

@DannyRomero19 wat grade r u in??

OpenStudy (anonymous):

now what is the height

OpenStudy (anonymous):

hey danny area of base isnt 8

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

what is formula of a square danny

OpenStudy (anonymous):

cummon dude!! 2*2 = 4 and 4*8 = 32!! u kiddin me or what?

OpenStudy (anonymous):

@DannyRomero19 that ws fr u

OpenStudy (anonymous):

@pavaneinstine what are u doing..... dont give direct answer things may be simple for u but it is difficult for others

OpenStudy (anonymous):

he got a point, but thnx anyways!! lol!!

OpenStudy (anonymous):

@sauravshakya cummon man! doesnt EVERYONE knw 2*2 = 4?? :O

OpenStudy (anonymous):

@pavaneinstine http://openstudy.com/study#/updates/501e9d76e4b0be43870ee34c

OpenStudy (anonymous):

idk why but i did 4*2 instead of 2*2

OpenStudy (anonymous):

PLZ help me with that problem @pavaneinstine